# Superconductivity in overdoped cuprates can be understood from a BCS perspective!

### v2

@@ I Introduction
Foreground: In this article (which is a perspective, and not a review), we analyze the evidence that the superconducting dome in an “ideal” disorder free cuprate would arise as a consequence of a crossover from a strongly correlated (and hence very complex) underdoped regime, where $T_{c}$ is intrinsically determined by phase ordering, to an overdoped regime, where a weak coupling perspective based on a Fermi liquid normal state and a BCS mean-field superconducting state is possible. As part of this analysis, we present reasons to attribute to the effects of disorder those observed features of actual overdoped cuprates that are inconsistent with a weak-coupling description. Our perspective is summarized in Figure 1, for which the essential assumptions are:
@@
The main point: Our primary purpose is to critically asses the validity of a weak coupling perspective as it relates to the overdoped cuprates. (While most theoretical studies of the cuprates have emphasized more exotic, strong coupling aspects of the problem, there is a significant subset that overlap either in some details or in perspective with the present approach [2, 3, 4, 5, 6, 7, 8, 9, 10, 11].) Specifically, we identify the experimental evidence that supports this perspective and, perhaps more importantly, observations that clearly contradict it. For example, several families of cuprates exhibit the “boomerang effect,” where $T_{\theta}$ decreases with increasing $p$ such that it remains comparable to $T_{c}$ throughout the overdoped regime[12, 13, 14, 15]. Moreover, when superconductivity is suppressed with a high magnetic field, “strange metal” behavior (as opposed to the expected Fermi-liquid behavior) is seen throughout the overdoped range of $p$ [16, 17]. We will discuss reasons to think that these features—those that challenge our perspective when taken at face value—are the consequence of intrinsic alloy-disorder.
@@
The perspective presented here is quite different from the widely discussed perspective in which optimal doping is associated with a quantum critical point whose influence dominates large portions of the phase diagram.[18, 19, 20] While our perspective does not preclude the existence of one or more quantum critical points with $p_{c}\lesssim p_{opt}$, one outcome of our discussion will be to adduce evidence—under the assumption that the weak-coupling analysis is reasonable for $p>p_{opt}$—that the pairing interaction does not support the idea that a putative quantum critical point at $p^{*}\sim p_{opt}$ plays a significant role in the pairing mechanism.
@@ IV Experiments on overdoped cuprates
There have been several theoretical proposals attempting to reconcile the unambiguous evidence from QO’s for the existence of a Fermi liquid with the observed transport anomalies. On purely phenomenological grounds, a “two fluid” model has been suggested,[48] although it is hard to see how such behavior could arise other than from some significant form of electronic inhomogeneity.[17] As discussed above, even a high-quality, “uniformly doped” sample will appear heterogeneous on short enough length scales due to the statistical distribution of dopant locations. This can lead to an electronically inhomogeneous state that consists of dilute superconducting “puddles” that might survive even to the high fields at which superconducting long-range coherence is destroyed; a recent paper[49] has demonstrated that scattering from the quantum phase fluctuations on such puddles can give rise to a $T$-linear resistivity. Alternatively, the possibility that anomalous scattering might be associated with local, nearly-quantum-critical patches of some other order - possibly associated with a disorder broadened antiferromagnetic quantum critical point - has been proposed in a series of related papers.[50, 51, 52, 53, 54]
@@
While in LSCO and Tl1201, the superfluid density (or more precisely $T_{\theta}$) decreases with increasing $p$ past optimal doping (the “boomerang effect” illustrated in the inset of Fig. 1), in YBa2Cu3Ox it increases monotonically as a function of $p$ across the entire accessible range of doping [56]. Particularly notable is the fact that $T_{\theta}$ increases by $50\%$ from its value at $p=0.16$, where $T_{\rm c}=94$ K, to $p=0.19$, where $T_{\rm c}=88$ K [57] (see Figure 2. This departs markedly from the Uemura relationship, where phase stiffness is proportional to the superconducting $T_{\rm c}$ [12].
@@
There are three classes of mechanisms that could suppress the macroscopic phase stiffness of overdoped cuprates relative to the bare Drude weight: i) intrinsic strong-coupling physics, ii) pair-breaking due to scattering, and iii) mesoscopic inhomogeneity (on scales of order the coherence length). All three may be present to some degree, but we argue that the latter is dominant. [^6] In Refs.[27, 28, 29], the authors argue that many salient features of the overdoped cuprates can be accounted for entirely on the basis of the pair-breaking effects of disorder, without considerations of disorder-induced inhomogeneous gap structures or phase fluctuations. While this analysis is consistent with our view of the “ideal” overdoped cuprate, it is not easily reconciled with the strong direct and indirect evidence of large amplitude gap inhomogeneities, nor with the fact that $T_{\theta}\sim T_{c}$ implies that phase fluctuations are significant.

### v3

@@ I Introduction
Foreground: In this article (which is a perspective, and not a review), we analyze the evidence that the superconducting dome in an “ideal” disorder free cuprate would arise from a crossover between a strongly correlated (and hence very complex) underdoped regime, where $T_{c}$ is intrinsically determined by phase ordering, to an overdoped regime, where it is possible to adopt a weak coupling “BCS” perspective based on a Fermi liquid normal state and a mean-field description of the superconducting state (but not, necessarily, invoking phonon-mediated pairing). As part of this analysis, we present reasons to attribute to the effects of disorder those observed features of actual overdoped cuprates that are inconsistent with a weak-coupling description. Our perspective is summarized in Figure 1, for which the essential assumptions are:
@@
The main point: Our primary purpose is to critically asses the validity of a weak coupling perspective as it relates to the overdoped cuprates. (While most theoretical studies of the cuprates have emphasized more exotic, strong coupling aspects of the problem, there is a significant subset that overlap either in some details or in perspective with the present approach [2, 3, 4, 5, 6, 7, 8, 9, 10, 11].) Specifically, we identify the experimental evidence that supports this perspective and, perhaps more importantly, observations that clearly contradict it. For example, several families of cuprates exhibit the “boomerang effect,” where $T_{\theta}$ decreases with increasing $p$ such that it remains comparable to $T_{c}$ throughout the overdoped regime (thus appearing to “turn around” when plotted parametrically as a function of $p$ on a $T_{\theta}$ vs $T_{\rm c}$ plot)[12, 13, 14, 15]. Moreover, when superconductivity is suppressed with a high magnetic field, “strange metal” behavior (as opposed to the expected Fermi-liquid behavior) is seen throughout the overdoped range of $p$ [16, 17]. We will discuss reasons to think that these features—those that challenge our perspective when taken at face value—are the consequence of intrinsic alloy-disorder.
@@
The conceptual underpinnings of the perspective presented here are distinct from those underlying the “two-fluid” phenomenology (2FP) discussed in Ayres et al. [18]. Whereas we identify as “extrinsic” (i.e. vanishing in the zero disorder limit) the various properties of overdoped materials that are inconsistent with a FL/BCS description, the 2FP interprets them as consequences of unspecified strong correlation effects that lead intrinsically to the existence of coexisting “coherent” and “incoherent” fluids of electronic excitations.
The present perspective is also fundamentally different from the widely discussed perspective in which optimal doping is associated with a quantum critical point whose influence dominates large portions of the phase diagram.[19, 20, 21] While our perspective does not preclude the existence of one or more quantum critical points with $p_{c}\lesssim p_{opt}$, one outcome of our discussion will be to adduce evidence—under the assumption that the weak-coupling analysis is reasonable for $p>p_{opt}$—that the pairing interaction does not support the idea that a putative quantum critical point at $p^{*}\sim p_{opt}$ plays a significant role in the pairing mechanism. (see the supplementary information for further discussion of the role of quantum criticality).
@@ III What would a BCS approach to the cuprates entail?
III.1 Additional considerations
In the Supplemental Materials, we expand upon certain important matters of perspective. In Sec. SM-I, we review the difference between a crossover and a phase transition in the context of the assumed crossover indicated by the rose region in Fig. 1. This is essential to the present analysis in that it justifies treating the mechanism of superconductivity in the overdoped regime without needing to account for the behavior of the underdoped regime, despite the fact that they are smoothly (adiabatically) connected to one another. In Sec. SM-II, we discuss aspects of the relation between the present perspective and perspectives that treat the existence of a putative quantum critical point at $p=p^{\star}$ under the superconducting dome as the central organizing feature of the cuprate phase diagram. In Sec. SM-III, we expand upon the tangible differences between homogeneous “dirty $d$-wave” and inhomogeneous, self-organized granularity as the cause of the suppressed superfluid density on the overdoped side of the phase diagram.
@@ IV Experiments on overdoped cuprates
There have been several theoretical proposals attempting to reconcile the unambiguous evidence from QO’s for the existence of a Fermi liquid with the observed transport anomalies. As discussed above, even a high-quality, “uniformly doped” sample will appear heterogeneous on short enough length scales due to the statistical distribution of dopant locations. This can lead to an electronically inhomogeneous state that consists of dilute superconducting “puddles” that might survive even to the high fields at which superconducting long-range coherence is destroyed; a recent paper[49] has demonstrated that scattering from the quantum phase fluctuations on such puddles can give rise to a $T$-linear resistivity. Experimental evidence for such a scenario comes from overdoped Tl2201, where increasing the magnetic field from 10 to 18 tesla reduces the $T$-linear slope by at least a factor of 2 [50]. Alternatively, the possibility that anomalous scattering might be associated with local, nearly-quantum-critical patches of some other order - possibly associated with a disorder broadened antiferromagnetic quantum critical point - has been proposed in a series of related papers.[51, 52, 53, 54, 55]
@@
While in LSCO and Tl1201, the superfluid density (or more precisely $T_{\theta}$) decreases with increasing $p$ past optimal doping (the “boomerang effect”), in YBa2Cu3Ox it increases monotonically as a function of $p$ across the entire accessible range of doping [57]. Particularly notable is the fact that $T_{\theta}$ increases by $50\%$ from its value at $p=0.16$, where $T_{\rm c}=94$ K, to $p=0.19$, where $T_{\rm c}=88$ K [58] (see Figure 2. This departs markedly from the Uemura relationship, where phase stiffness is proportional to the superconducting $T_{\rm c}$ [12].
@@
There are three classes of mechanisms that could suppress the macroscopic phase stiffness of overdoped cuprates relative to the bare Drude weight: i) strong-coupling physics, ii) pair-breaking due to scattering,[28, 29, 30] and iii) mesoscopic inhomogeneity (on scales of order the coherence length). The first mechanism is intrinsic, while the later two are extrinsic in the sense that they disappear in the limit of vanishing disorder. While all three effects may be present to some degree, we argue that extrinsic effects are dominant. Further, we suggest that mesoscopic inhomogeneity provides a more compelling explanation for several experimental phenomena, although it should be noted that the debate over the relative importance of pair breaking versus inhomogenaity is still ongoing [28, 29, 30]. We discuss this issue further in the supplementary information.